### v1

@@ Abstract.
bimagma, dependence, function semiring, hyperfield, hyperpair, Krasner, linear algebra, magma, matrix, metatangible, pair, Property N, polynomial, residue semiring, subgroup, semiring, quotient semiring, supertropical, root, tensor, vector space.
@@ 1. Overview
Semirings can be studied as $\Omega$-algebras in universal algebra [21, Chapter 2]. In particular, one can define the direct product of semirings $\prod_{i\in I}\mathcal{A}_{i}.$ One calls two structures isomorphic (written $\cong$) when there is a 1:1 correspondence which respects their operations and distinguished elements (in the case of semirings, $\cdot,\mathbf{0},\mathbf{1}$).
@@ 1.7.5. Monoid semirings
The polynomial semiring We write $\mathcal{A}[\Lambda]$ for the semiring $\mathcal{A}[\mathcal{M}]$, when $\mathcal{M}$ is the word monoid in $t$ commuting letters $\lambda_{1},\dots,\lambda_{t}$, and denoted $\mathcal{A}[\lambda]$ when $t=1.$ The semiring $\mathbb{N}[\lambda]$ plays a key role in the work of Lorscheid [38, 39] in $\mathbb{F}_{1}$-geometry.
@@ 1.8. Modules and semirings over a monoid
Just as one studies rings as algebras over an underlying commutative ring (often a field), we should like to study semirings over an underlying commutative semiring, but it turns out that the key is its multiplicative structure, which is a monoid.
@@
    We shall assume henceforth that $\mathcal{A}$ is a faithful $\mathcal{T}$-module, in the sense that if $ab_{1}=ab_{2}$, for $a\in\mathcal{T}$ and $b_{1},b_{2}\in\mathcal{A}$, then $b_{1}=b_{2}.$ Our convention is to write $a$ for an element of $\mathcal{T},$ and $b$ for an element of $\mathcal{A}.$ We call $\mathcal{T}$ the underlying monoid of tangible elements.
@@ A $\mathcal{T}_{1}$-homomorphism $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a map satisfying
  • As explained in [21, Chapter 2], in universal algebra, homomorphisms are defined via congruences, where we recall that a congruence on $\mathcal{A}$ is an equivalence relation which viewed as a subset of $\mathcal{A}\times\mathcal{A}$ is a subalgebra. In particular, the trival congruence $\operatorname{diag}$ is $\{(a,a):a\in\mathcal{A}\}$. This approach fits in well with semigroups and semirings viewed via universal algebra, but does not always yield the categories that we desire, as we shall see, and other morphisms are described in §8.
@@
  • If $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a homomorphism then for $b\in\mathcal{A}_{2},$ the inverse image $f^{-1}(b)$ is an ideal, not a congruence. This creates a certain tension, where sometimes we choose ideals and sometimes congruences, to be explained in Note 2.24.
@@ 2. Pairs
    From now on, excepting the Examples (iv),(v) of §2.1.3 and Definition mapr, the underlying monoid $\mathcal{T}$ is given ( until §8.1), and we simply call a $\mathcal{T}$-pair a “pair.” We call $\mathcal{A}_{0}$ the “null part” of $\mathcal{A}.$
@@ 2.1. Semiring Pairs
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be
    • proper if $\mathcal{A}_{0}\cap\mathcal{T}_{0}=\{\mathbf{0}\}$.
    • shallow if $\mathcal{T}\cup\mathcal{A}_{0}=\mathcal{A}.$
  3. (3)
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is of the first kind if $\mathbf{1}+\mathbf{1}\in\mathcal{A}_{0}$ and of the second kind if $\mathbf{1}+\mathbf{1}\notin\mathcal{A}_{0}$ (And thus $a+a\notin\mathcal{A}_{0}$ for all $a\in\mathcal{T}$).
  4. (4)
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is an $\mathcal{A}_{0}$-domain if $\mathcal{A}_{0}$ is a strongly prime ideal, i.e., $b_{1}b_{2}\in\mathcal{A}_{0}$ implies $b_{1}\in\mathcal{A}_{0}$ or $b_{2}\in\mathcal{A}_{0}$.
MAJOR NOTE 2.15.
“Classical algebra” is when $R$ is an algebra over a commutative ring $C$, and $\mathcal{T}_{0}=C$, viewed as a multiplicative monoid. Or we could simply take $\mathcal{T}=\{\mathbf{1}\}.$ But in this paper we are more interested in other situations.
The idea here is that in proper semiring pairs we “see” the elements $\mathcal{T},$ but $\mathcal{A}_{0}$ controls much of the structure. Historically, “first kind” is important since it gives Izhakian’s supertropical algebra “supertropical algebra” and the “flavor” of characteristic 2. “Second kind” is more closely related to classical algebra of characteristic $\neq 2$.
@@
2.1.1. Property N
A key issue is how $\mathcal{T}$ interfaces with $\mathcal{A}.$
Definition 2.17.
  1. (i)
    $e:=\mathbf{1}+\mathbf{1}^{\dagger}\in\mathcal{A}_{0}$.
  2. (ii)
  3. (iii)
    $b^{\circ}\in\mathcal{A}_{0}$ for each $b\in\mathcal{A}$.
Lemma 2.18 ([4, Lemma 3.10]).
$\mathbf{1}^{\dagger}e=e.$ Thus, in a semiring pair, $e^{2}=e+e.$
MAJOR NOTE 2.19.
  1. (i)
  2. (ii)
    In earlier incarnations, $\mathbf{1}^{\dagger}$ was required to be unique, designated $(-)\mathbf{1}$, with $((-)\mathbf{1})^{2}=\mathbf{1},$ yielding a unique “negation map” $b\mapsto((-)\mathbf{1})b.$ In this case we write ()b(-)b for $b^{\dagger},$ and say that the pair $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated. Then we write $b_{1}(-)b_{2}$ for $b_{1}+((-)b_{2}).$
Examples are given in [4, §5], and below in §2.1.3.
@@ 2.1.3. Examples of pairs
    Some miscellaneous examples:
    • $\mathbb{V}_{n}$ example 1.3 becomes a pair satisfying Property N, when we put $\mathcal{T}=\{1\}$ and $\mathcal{A}_{0}=\{\mathbf{0},n\}.$
@@
    • Suppose $\mathcal{A}=\mathbb{N}[\Lambda]$ in commuting indeterminates $\{\lambda_{i}:1\leq i<\infty\}$ over the monoid $\mathcal{T}$ generated by the $\lambda_{i},$ We identify all elements of $\mathcal{A}$ of support at least of order 3 with $e$. $(\mathcal{A},\mathcal{A}_{0})$ is not weakly metatangible since $\lambda_{1}+\lambda_{2}\notin\mathcal{T}+\mathcal{A}_{0}.$
@@
    Any $\mathcal{T}$-semiring $\mathcal{A}$ gives rise to a doubled pair where $\hat{A}$ is as in Definition 1.18, and the null part $\hat{\mathcal{A}}_{0}$ is as in Remark 1.19. $(\hat{\mathcal{A}},\hat{\mathcal{A}}_{0})$ has the negation map $(-)(b_{1},b_{2})=(b_{2},b_{1}).$
@@
      The monoid semiring pair. If $\mathcal{M}$ is a monoid then $(\mathcal{A}[\mathcal{M}],\mathcal{A}_{0}[\mathcal{M}])$ is a semiring pair over the monoid $\mathcal{T}\cdot\mathcal{M}=\{ay:a\in\mathcal{T},\ y\in\mathcal{M}\}$.
    In each case of (v), we have altered $\mathcal{T};$ this will be discussed more formally in §8.1. But then, even if $(\mathcal{A},\mathcal{A}_{0})$ is metatangible, $(\mathcal{A}^{S},\mathcal{A}_{0}^{S})$ and $(\mathcal{A}[S],\mathcal{A}_{0}[S])$ need not be weakly metatangible.
2.1.4. Hyperpairs
@@ 2.2. Distributivity
    The analogous definitions from 2.14 hold for nd-semiring pairs.
@@ 3. Roots of polynomials
If $b$ is a root of $f,$ then the paired homomorphism $(\mathcal{A},\mathcal{A}_{0})[\lambda]\to(\mathcal{A},\mathcal{A}_{0})[b]$ sending $\lambda\mapsto b$ sends $f$ to an element of $\mathcal{A}_{0}.$
@@ 3.1. Factorization with respect to roots of polynomials
  • Usual division of elements in a monoid: $f_{1}|f_{2},$ if there is tangible $g$ such that $f_{2}=gf_{1}.$
@@
Gunn [Proposition 4.7][17] has an argument for roots being $\preceq$-roots in a certain instance. We modify his argument to provide a rather general fundamental result.
@@
    If $\preceq$ satisfies fissure, then each tangible root of any tangible polynomial is a $\preceq$-root.
Proof.   (i) Suppose $f\,|_{\preceq}\,(\lambda(-)a)g=\lambda g(-)ag.$ Write $f=\sum\alpha_{i}\lambda^{i}$ and $g=\sum\beta_{i}\lambda^{i}.$ Matching coefficients we get $\alpha_{i}\preceq(-)a\beta_{i}+\beta_{i-1}$, $i\geq 1,$ implying $(-)\beta_{i-1}\preceq(-)\alpha_{i}(-)a\beta_{i}$ since $\beta_{i-1}$ is presumed tangible, so $\beta_{i-1}\preceq\alpha_{i}+a\beta_{i}$, and $\alpha_{0}\preceq(-)a\beta_{0},$ implying $\mathbf{0}\preceq\alpha_{0}+a\beta_{0}\preceq\alpha_{0}+a\alpha_{1}+a^{2}\beta_{1}\preceq\sum_{i=0}^{n-1}\alpha_{i}a^{i}+a^{n}\beta_{n}=\sum_{i=0}^{n-1}\alpha_{i}a^{i}+\alpha_{n}a^{n}=f(a).$
@@
Proof.   (i) $(\Rightarrow)$ (ii) by Theorem 3.13(ii) Lemma 3.9 and (iii).
@@
We can treat several $\preceq$-roots at once if $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}$-prime.
@@
If $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}$-prime, then a monic polynomial $f$ cannot have more than $n$ distinct tangible $\preceq$-roots. If $a_{1},\dots,a_{n}$ are distinct tangible $\preceq$-roots then $f\preceq\prod_{i=1}^{n}(\lambda(-)a_{i}).$
Proof.   Writing $f\preceq(\lambda(-)a_{1})g$ for $g$ monic, we have $\mathbf{0}\leq f(a_{2})=(a_{2}(-)a_{1})g(a_{2}).$ By hypothesis, $g(a_{2})\in\mathcal{A}_{0},$ clearly of degree n1.n-1. Continuing inductively, and noting $\deg g=\deg f-1,$ we cannot have more than $n$ tangible $\preceq$-roots of $f$.
Corollary 3.16.
The Hamilton-Cayley polynomial of a matrix cannot have more than $n$ $\preceq$-roots.
@@
    (Like Gunn’s version) $a$ is a double $\preceq$-root of $f$ if, for some tangible polynomial $g$, $f(-)(\lambda(-)a)g\in\mathcal{A}_{0}$ and $a$ is a $\preceq$-root of $g$.
@@
These three versions are known to be equivalent in classical algebra, so which should we choose? Version (i) may seem the most natural at this stage, but we would like a double root also to be a single root. If $(\lambda(-)a)^{2}h\succeq f$ then clearly $(\lambda(-)a)g\succeq f$ for $g=(\lambda(-)a)h,$ but unfortunately $(\lambda(-)a)h$ need not be tangible. This leads to (ii), in which Descartes’ rule of signs has been proved for tropical extensions, in [6, 17, 5], but there is ambiguity with respect to $g$, which leads to complications. Therefore we turn to (iii), which is straightforward and unambiguous, but unfortunately $f^{\prime}$ need not be tangible! But $f^{\prime}$ indeed is tangible when no multiple of $a$ is in $\mathcal{A}_{0}$.

### v2

@@ Abstract.
dependence, function semiring, hyperfield, hyperpair, Krasner, linear algebra, magma, matrix, metatangible, pair, Property N, polynomial, residue semiring, subgroup, semiring, quotient hyperring, supertropical, root, tensor, vector space.
@@ 1. Overview
Semirings can be studied as $\Omega$-algebras in universal algebra [28, Chapter 2]. In particular, one can define the direct product of semirings $\prod_{i\in I}\mathcal{A}_{i}.$ One calls two structures isomorphic (written $\cong$) when there is a 1:1 correspondence which respects their operations and distinguished elements (in the case of semirings, $\cdot,\mathbf{0},\mathbf{1}$).
@@ 1.7.5. Monoid semirings
The polynomial semiring. We write $\mathcal{A}[\Lambda]$ for the semiring $\mathcal{A}[\mathcal{M}]$, when $\mathcal{M}$ is the word monoid in $t$ commuting letters $\lambda_{1},\dots,\lambda_{t}$, and denoted $\mathcal{A}[\lambda]$ when $t=1.$ The semiring $\mathbb{N}[\lambda]$ plays a key role in the work of Lorscheid [38, 39] in $\mathbb{F}_{1}$-geometry.
@@ 1.8. Modules and semirings over a monoid
Just as one studies rings as algebras over an underlying commutative ring (often a field), we might be inclined to study semirings over an underlying commutative semiring, but it turns out that the key is its multiplicative structure, which is a monoid.
@@
    We call $\mathcal{T}$ the underlying monoid of tangible elements. Our convention is to write $a$ for an element of $\mathcal{T},$ and $b$ for an element of $\mathcal{A}.$ We shall assume henceforth that $\mathcal{A}$ is a faithful $\mathcal{T}$-module, in the sense that if $ab_{1}=ab_{2}$, for $a\in\mathcal{T}$ and $b_{1},b_{2}\in\mathcal{A}$, then $b_{1}=b_{2}.$
@@ A $\mathcal{T}_{1}$-homomorphism $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a map satisfying
  • As explained in [28, Chapter 2], in universal algebra, homomorphisms are defined via congruences, where we recall that a congruence on $\mathcal{A}$ is an equivalence relation which viewed as a subset of $\mathcal{A}\times\mathcal{A}$ is a subalgebra. In particular, the trival congruence $\operatorname{diag}$ is $\{(a,a):a\in\mathcal{A}\}$. This approach fits in well with semigroups and semirings viewed via universal algebra, but does not always yield the categories that we desire, as we shall see, and other morphisms are described in §8.
@@
  • If $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a homomorphism then for $b\in\mathcal{A}_{2},$ the inverse image $f^{-1}(b)$ is an ideal, not a congruence. This creates a certain tension, where sometimes we choose ideals and sometimes congruences, to be explained in Note 2.25.
@@ 2. Pairs
    From now on, excepting the Examples (iii), (iv), and (v) of §2.1.2 and Definition 5.1, the underlying monoid $\mathcal{T}$ is given (until §8.1), and we simply call a $\mathcal{T}$-pair a “pair.” We call $\mathcal{A}_{0}$ the “null part” of $\mathcal{A}.$
@@
Definition 2.3.
  1. (1)
    A pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be
    • proper if $\mathcal{A}_{0}\cap\mathcal{T}_{0}=\{\mathbf{0}\}$.
    • shallow if $\mathcal{T}\cup\mathcal{A}_{0}=\mathcal{A}.$
  2. (2)
    A pair $(\mathcal{A},\mathcal{A}_{0})$ is of the first kind if $\mathbf{1}+\mathbf{1}\in\mathcal{A}_{0}$ and of the second kind if $\mathbf{1}+\mathbf{1}\notin\mathcal{A}_{0}$ (And thus $a+a\notin\mathcal{A}_{0}$ for all $a\in\mathcal{T}$).
2.0.1. Property N
A key issue is how $\mathcal{T}$ interfaces with $\mathcal{A}.$
Definition 2.4.
  1. (i)
    $e:=\mathbf{1}+\mathbf{1}^{\dagger}\in\mathcal{A}_{0}$.
  2. (ii)
  3. (iii)
    $b^{\circ}\in\mathcal{A}_{0}$ for each $b\in\mathcal{A}$.
Lemma 2.5 ([4, Lemma 3.10]).
$\mathbf{1}^{\dagger}e=e.$ Thus, in a semiring pair, $e^{2}=e+e.$
MAJOR NOTE 2.6.
  1. (i)
  2. (ii)
    In earlier incarnations, $(\mathbf{1}^{\dagger})$ was designated $(-)\mathbf{1}$, with the extra requirement $((-)\mathbf{1})^{2}=\mathbf{1},$ yielding a “negation map” $b\mapsto((-)\mathbf{1})b.$ In this case we write ()b(-)b for $b^{\dagger},$ and we write $b_{1}(-)b_{2}$ for $b_{1}+((-)b_{2}).$ The pair $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated if $(-)\mathbf{1}$ is unique.
Examples are given in [4, §5], and below in §2.1.2.
@@ 2.1. Semiring Pairs
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is an $\mathcal{A}_{0}$-domain if $b_{1}b_{2}\in\mathcal{A}_{0}$ implies $b_{1}\in\mathcal{A}_{0}$ or $b_{2}\in\mathcal{A}_{0}$.
MAJOR NOTE 2.19.
“Classical algebra” is when $\mathcal{A}$ is an algebra over a commutative ring $C$; $\mathcal{T}_{0}=C$, viewed as a multiplicative monoid, and $\mathcal{A}_{0}=\{0\}$ Or we could simply take $\mathcal{T}=\{\mathbf{1}\}.$
But in this paper we are more interested in other situations. The idea is that in proper semiring pairs we “see” the elements $\mathcal{T},$ but $\mathcal{A}_{0}$ controls much of the structure. Historically, “first kind” is important since it gives Izhakian’s supertropical algebra “supertropical algebra” and the “flavor” of characteristic 2. “Second kind” is more closely related to classical algebra of characteristic $\neq 2$.
@@ 2.1.2. Examples of semiring pairs
    Some miscellaneous examples, taken from [4]:
    • $\mathcal{A}=\mathbb{V}_{n}$ (Example 1.3) becomes a pair of first kind satisfying Property N, when we put $\mathcal{T}=\{1\}$ and $\mathcal{A}_{0}=(\mathcal{A}\cap 2\mathbb{N})\cup\{n\}.$
@@
    • Suppose $\mathcal{A}=\mathbb{N}[\Lambda]$ in commuting indeterminates $\{\lambda_{i}:1\leq i<\infty\}$ with underlying monoid $\mathcal{T}$ generated by the $\lambda_{i},$ We identify all elements of $\mathcal{A}$ of support at least of order 3 with $e$. $(\mathcal{A},\mathcal{A}_{0})$ is not weakly metatangible since $\lambda_{1}+\lambda_{2}\notin\mathcal{T}+\mathcal{A}_{0}.$
@@
    Any $\mathcal{T}$-semiring $\mathcal{A}$ gives rise to a doubled pair with underlying monoid $(\mathcal{T}\times\mathbf{0})\cup(\mathbf{0}\times\mathcal{T}),$ where $\hat{A}$ is as in Definition 1.18, and the null part $\hat{\mathcal{A}}_{0}$ is as in Remark 1.19. $(\hat{\mathcal{A}},\hat{\mathcal{A}}_{0})$ has the negation map $(-)(b_{1},b_{2})=(b_{2},b_{1}).$
@@
      The monoid semiring pair. If $\mathcal{M}$ is a monoid then $(\mathcal{A}[\mathcal{M}],\mathcal{A}_{0}[\mathcal{M}])$ is a semiring pair over with underlying monoid $\mathcal{T}\cdot\mathcal{M}=\{ay:a\in\mathcal{T},\ y\in\mathcal{M}\}$.
    In (iii), (iv), and (v), we have altered $\mathcal{T};$ this will be discussed more formally in §8.1. But then, even if $(\mathcal{A},\mathcal{A}_{0})$ is metatangible, $(\mathcal{A}^{S},\mathcal{A}_{0}^{S})$ and $(\mathcal{A}[S],\mathcal{A}_{0}[S])$ need not be weakly metatangible.
2.2. Hyperpairs
@@ 2.3. Distributivity
    The analogous definitions from 2.18 hold for nd-semiring pairs.
@@
  2. (ii)
    When the $\mathcal{H}$ is stringent, the hyperpair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ is weakly metatangible and has a shallow sub-hyperpair $(\mathcal{P}(\mathcal{H})+\mathcal{H},,\mathcal{P}(\mathcal{H})_{0})$.
@@ 3. Roots of polynomials
We call a polynomial tangible if its coefficients are in $\mathcal{T},$ despite the fact that the product of two polynomials need not be tangible. Now, if $f$ is tangible and ${\mathbf{a}}\in\mathcal{T}^{(I)},$ then
@@
The map $\mathcal{A}[\Lambda]\to\mathcal{A}$ given by $\Lambda\mapsto{\mathbf{a}}$ is a homomorphism if ${\mathbf{a}}\in\mathcal{T},$ which yields a paired homomorphism $(\mathcal{A},\mathcal{A}_{0})[\Lambda]\to(\mathcal{A},\mathcal{A}_{0})$, sending $f$ to an element of $\mathcal{A}_{0}$ if and only if $\mathbf{a}$} is a root of $f.$
@@ 3.1. Factorization with respect to roots of polynomials
  • Usual division of elements in a monoid: $f_{1}|f_{2},$ if there is a tangible polynomial $g$ such that $f_{2}=gf_{1}.$
@@
Gunn [17, Proposition 4.7] has an argument for roots being $\preceq$-roots in a certain instance. We modify his argument to provide a rather general fundamental result.
@@
    If $\preceq$ satisfies fissure, then every tangible root of each tangible polynomial is a $\preceq$-root.
Proof.   (i) Suppose $f\,|_{\preceq}\,(\lambda(-)a)g=\lambda g(-)ag.$ Write $f=\sum\alpha_{i}\lambda^{i}$ and $g=\sum\beta_{i}\lambda^{i}.$ Matching coefficients we get $\alpha_{i}\preceq(-)a\beta_{i}+\beta_{i-1}$, $i\geq 1,$ implying $(-)\beta_{i-1}\preceq(-)\alpha_{i}(-)a\beta_{i}$ since $\beta_{i-1}$ is presumed tangible, so $\beta_{i-1}\preceq\alpha_{i}+a\beta_{i}$, and $\alpha_{0}\preceq(-)a\beta_{0},$ implying
$\mathbf{0}\preceq\alpha_{0}+a\beta_{0}\preceq\alpha_{0}+a\alpha_{1}+a^{2}\beta_{1}\preceq\sum_{i=0}^{n-1}\alpha_{i}a^{i}+a^{n}\beta_{n}=\sum_{i=0}^{n-1}\alpha_{i}a^{i}+\alpha_{n}a^{n}=f(a).$
@@
Proof.   (i) $(\Rightarrow)$ (ii) by Theorem 3.13(ii) and Lemma 3.9(iii).
@@
We often can treat several $\preceq$-roots simultaneously if $(\mathcal{A},\mathcal{A}_{0})$ is a $\mathcal{A}_{0}$-domain.
@@
Suppose that a pair $(\mathcal{A},\mathcal{A}_{0})$ is a $\mathcal{A}_{0}$-domain with unique negation, such that either $\mathcal{A}$ is a semiring or $(\mathcal{A},\mathcal{A}_{0})$ is shallow. Then a monic polynomial $f$ cannot have more than $n$ distinct tangible $\preceq$-roots. If $a_{1},\dots,a_{n}$ are distinct tangible $\preceq$-roots then $f\preceq\prod_{i=1}^{n}(\lambda(-)a_{i}).$
Proof.   Writing $f\preceq(\lambda(-)a_{1})g$ for $g$ monic, we have $\mathbf{0}\leq f(a_{2})=a_{2}g(a_{2})(-)a_{1}g(a_{2}).$
If $\mathcal{A}$ is a semiring then $(a_{1}(-)a_{2})g(a_{2})\in\mathcal{A}_{0},$ implying $g(a_{2})\in\mathcal{A}_{0}.$
If $(\mathcal{A},\mathcal{A}_{0})$ is shallow and $g(a_{2})\in\mathcal{T}$, then $a_{2}g(a_{2})(-)a_{1}g(a_{2})=(a_{2}(-)a_{1})g(a_{2}),$ implying $a_{2}(-)a_{1}\in\mathcal{A}_{0},$ so $a_{1}=a_{2}.$ Hence $g(a_{2})\in\mathcal{A}_{0},$ clearly of degree n1.n-1. Continuing inductively, and noting $\deg g=\deg f-1,$ we cannot have more than $n$ tangible $\preceq$-roots of $f$.
For application to hyperfields, see Remark 2.31.
@@
    (Like Gunn’s version) $a$ is a double $\preceq$-root of $f$ if, for some tangible polynomial $g$, $f(-)(\lambda(-)a)g\in\mathcal{A}_{0}[\lambda]$ and $a$ is a $\preceq$-root of $g$.
@@
    $a$ is a double $\preceq$-root of $f$ if, for some tangible polynomial $g$, $f(-)(\lambda(-)a\preceq g$ and $a$ is a $\preceq$-root of $g$.
  4. (iv)
@@
These four versions are known to be equivalent in classical algebra, so which should we choose? Version (i) may seem the most natural at this stage, but we would like a double root also to be a single root. If $(\lambda(-)a)^{2}h\succeq f$ then clearly $(\lambda(-)a)g\succeq f$ for $g=(\lambda(-)a)h,$ but unfortunately $(\lambda(-)a)h$ need not be tangible. This leads to (ii), in which Descartes’ rule of signs has been proved for tropical extensions, in [6, 17, 5], but there is ambiguity with respect to $g$, which leads to complications. We could turn to (iv), which is straightforward and unambiguous, but unfortunately $f^{\prime}$ need not be tangible! But $f^{\prime}$ indeed is tangible when no multiple of $a$ is in $\mathcal{A}_{0}$. We favor (iii), which is closest to a repetition of Definition 3.12, and which we intend to use in future work.